OpenStudy (anonymous):

Find an equation of the line through the midpoint (-7,-4) and (4,14) perpendicular to the line 2x-4y=5

OpenStudy (anonymous):

original line: y=1/2-1.25 midpoint: (-1.5,5) use the y-y1=m(x-x1) m is ypur slope from the original line but the reciprocal which is 2: i think the answer is y=2x+8